I need help solving an Algebra problem concerning Rational Equations.
Here is the equation.
P=Po/t+1= P=4000/t+1 =
If a town has a population of 4000 how many years would it take the
towns population to decrease to 600?

Answer by Alan3354(69443)

P=Po/t+1= P=4000/t+1 =
If a town has a population of 4000 how many years would it take the
towns population to decrease to 600?
------------
600 = 4000/(t+1)
t+1 = 4000/600 = 20/3
3(t+1) = 20
3t+3 = 20
3t = 17
t = 17/3 years